### v2

# Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach

Abstract
The Lucas critique emphasizes the importance of considering the impact of policy changes on the expectations of micro-level agents in macroeconomic policymaking. However, the inherently self-interested nature of large-scale micro-agents, who pursue long-term benefits, complicates the formulation of optimal macroeconomic policies. This paper proposes a novel general framework named Dynamic Stackelberg Mean Field Games (Dynamic SMFG) to model such policymaking within sequential decision-making processes, with the government as the leader and households as dynamic followers. Dynamic SMFGs capture the dynamic interactions among large-scale households and their response to macroeconomic policy changes. To solve dynamic SMFGs, we propose the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm, which leverages the population distribution of followers to represent high-dimensional joint state and action spaces. In experiments, our method surpasses macroeconomic policies in the real world, existing AI-based and economic methods. It allows the leader to approach the social optimum with the highest performance, while large-scale followers converge toward their best response to the leader’s policy. Besides, we demonstrate that our approach retains effectiveness even when some households do not adopt the SMFG policy. In summary, this paper contributes to the field of AI for economics by offering an effective tool for modeling and solving macroeconomic policy-making issues.
1 Introduction
Macroeconomic policymaking is fundamental to the sustained development of an economy [52, 46]. Governments influence economic production, wealth distribution, social stability, and welfare through economic policies such as interest rates, taxation, and fiscal spending [49]. Therefore, accurately modeling and solving optimal macroeconomic policies and simulating the effects of their implementation are critical issues. According to the Lucas critique [39], predicting the impact of macroeconomic policies effectively requires modeling individual micro-level behavior—termed "microfoundations"—to forecast personal responses to policy changes and aggregate these into macroeconomic effects. Nonetheless, several challenges persist: dynamic interactions among micro-agents, between micro-agents and macro-authorities, the inherently selfish behavior of agents focused on optimizing long-term returns, and the curse of dimensionality due to the large-scale nature of agents such as households and firms.
This paper introduces a novel framework termed Dynamic Stackelberg Mean Field Game (Dynamic SMFG) designed to address complex macroeconomic policy-making problems. Within this dynamic setting, the government, serving as the leader, adjusts policies iteratively, while a large population of households, acting as followers, respond accordingly. This process involves sequential decision-making, with both the leader and the followers aiming to optimize their cumulative long-term benefits. Given the leader’s policy, the SMFG reduces to a mean field game, which is widely used for modeling interactions among large-scale homogenous agents [33]. In Dynamic SMFGs, followers aim to maximize their long-term benefits by providing the best response to the leader’s policy, while the leader’s goal is to find an optimal policy that effectively guides the massive followers toward achieving comprehensive macroeconomic objectives. Existing approaches often simplify Dynamic SMFGs to single-step decisions [45] or assume myopic followers [63], which do not adequately capture the complexities of macroeconomic policy-making. Therefore, we propose a data-driven reinforcement learning algorithm, Stackelberg Mean-Field Reinforcement Learning (SMFRL), to solve Dynamic SMFGs without requiring transition information or prior knowledge. The SMFRL algorithm utilizes a mean field approximation [28] to reduce the high-dimensional joint state and actions of followers to a manageable population distribution, allowing for the effective learning of a central critic Q function based on the overall followers’ distribution.
Experiments conducted on TaxAI [41], an environment designed for studying optimal tax policies, first validate the performance of SMFRL in solving Dynamic SMFGs. The results indicate that SMFRL enables followers to closely achieve their best response while allowing the leader to approach the social optimum with the highest performance compared with baselines. Furthermore, the paper compares the Dynamic SMFG method with macroeconomic policies such as AI Economist [62], the Saez tax [50], the 2022 U.S. Federal tax, and free market policy [5], demonstrating the superiority of the Dynamic SMFG framework in macroeconomic policymaking. Considering real-world applications, our experiments show that even if a portion of households do not adopt the follower policy learned by the SMFG method, it does not adversely affect the utility, wealth, and income of households that do adopt the SMFG policy, which are in fact higher. However, the GDP of the economy may decrease with an increase in the number of households not supporting the SMFG policy, as they typically contribute less labor. In summary, the contributions of this paper are threefold:
1. It innovatively proposes a general framework based on Dynamic Stackelberg Mean Field Games to model macroeconomic policy-making, considering key issues such as microfoundations, dynamic interactions and feedback, rational expectations, and randomness.
2. It introduces a data-driven Stackelberg Mean Field Reinforcement Learning algorithm to solve dynamic SMFGs without environmental transition information and prior knowledge, validated experimentally for its effectiveness.
3. It empirically demonstrates the superiority of the SMFG method in modeling and solving optimal macroeconomic policy problems and discusses its robustness in real-world scenarios.
The source code for the SMFG method is available in an anonymous GitHub repository  [^1] https://anonymous.4open.science/r/SMFG_macro_7740.
2 Related work
2.1 Economic Methods
In the field of economics, research on economic policy typically relies on theoretical model studies, empirical analysis, and econometric methods. Classic theoretical models, such as the IS-LM model [24, 22], are applicable for analyzing short-term policy effects but overlook long-term factors and price fluctuations. The aggregate demand-aggregate supply (AD-AS) model [19, 34] focuses on the relationship between aggregate demand and supply, integrating short-term and long-term factors, but simplifies macroeconomic dynamics. The Solow growth model [10] provides a framework for economic growth but does not consider market imperfections and externalities. New Keynesian models [9, 21] emphasize the stickiness of prices and wages, suitable for explaining economic fluctuations and policy interventions, but are based on strong assumptions. DSGE models [1, 55], built on micro-foundations for macroeconomic forecasting, offer a consistent framework for policy analysis but are limited in handling nonlinearities and market imperfections. Empirical analysis [47, 59] primarily evaluates policy effects through historical data but faces limitations in data timeliness and accuracy. Econometric methods [30, 15], such as regression and time series analysis, can quantify policy effects but are constrained by model settings and data quality.
2.2 AI for Economics
Artificial intelligence (AI), such as reinforcement learning, may offer new perspectives for solving complex economic problems [56]. In terms of macroeconomics, AI Economist [62] employs curriculum learning on tax policy design, [58] studies the collaboration between central and local governments under COVID-19, as well as monetary policy [26, 12], international trade [51], market pricing [14]. However, these studies consider the government as a single agent, neglecting the dynamic response of the households to policies. [31] studies democratic AI by human voting models, [60] examines the impact of social sanction rules on labor division. However, these works are based on simplified settings and remain a gap in real-world economic policy. In terms of microeconomics, [53, 48, 3] study the optimal saving and consumption problems of micro-agents, and other researchers explore rational expectation equilibrium [32, 25], multiple equilibria under real-business-cycle [13], the emergence of barter behavior [29], optimal asset allocation and savings strategies [44]. In summary, research on macroeconomic problems still lacks a bridge framework of macro policy and microfoundations, which has drawn our attention to the Stackelberg mean field game.
Figure 1: The mapping from optimal macroeconomic policy problems (left) to the Dynamic Stackelberg Mean Field Games (right). The left represents the reality where the macro-level government implements fiscal or monetary policies to optimize GDP, welfare, or equity, and the micro-level consists of large-scale households engaging in economic activities. In dynamic SMFGs, the government acts as the leader, and households as followers, respond to the leader’s policy.
2.3 Stackelberg Mean Field Game
Stackelberg Mean Field Games (SMFG) model interactions between the leader and numerous homogeneous followers, employing both model-based and model-free approaches. Model-based methods, such as those presented in [23], treat SMFGs as minimax optimization problems, streamlining the leader’s strategy to a single optimal action rather than a comprehensive policy. Studies by [7, 20, 16, 8] utilize forward-backward stochastic functional differential equations to determine the followers’ mean field equilibrium before computing the leader’s optimal strategy. This includes addressing terminal state constraints [20] and simplifying scenarios via linear-quadratic formulations [6, 37, 43, 18, 27]. Such methods, though theoretically robust, often rely on simplifying real-world complexities and require extensive knowledge of environmental transitions. Extending SMFG to model epidemic dynamics [4] showcases its broader applicability in public health management. Conversely, model-free approaches leverage complex, data-rich environments without the need for prior environmental knowledge. [11] employs a modified deep backward stochastic differential equation framework to establish followers’ equilibrium, despite constraints on leader decisions. [45] investigates applications in privacy protection through a single-step decision model. [42] delves into locally optimal defensive strategies given predefined attacker trajectories. [35] learns followers’ transition functions from empirical data, and facilitates SMFG solutions through the empirical Fokker-Planck equation.
However, in real-world settings, the sequential decision-making between a leader and numerous followers to optimize long-term gains is common but inadequately studied. This paper proposes a data-driven reinforcement learning algorithm for solving the aforementioned problem, termed the Dynamic Stackelberg Mean Field Game, without reliance on environmental transition information or prior knowledge.
3 Stackelberg Mean Field Game Modeling
3.1 Macroeconomic Policy-Making
Macroeconomic policy-making involves formulating effective policies to regulate and guide an economy toward achieving sustainable growth, low unemployment, and stable prices. In addressing these challenges, several key elements must be considered: ① Microfoundations: According to Lucas critique, basing macroeconomic models on microeconomic behaviors is crucial. This ensures that the actions of individual agents, such as households and firms, aggregate to produce macroeconomic outcomes. ② Dynamic Interactions: Extensive interactions occur both between the macro-agent (the government) and a multitude of micro-agents (households), and within groups of households. These interactions can lead to complex economic dynamics and endogenous shocks. ③ Rational Expectations: Agents are typically modeled as rational and self-interested, making decisions based on available information and their expectations about future policies and economic conditions. ④ Stochastic Elements: Incorporating randomness and shocks, such as technological changes, policy shifts, and external economic shocks, is essential for understanding variability in economic outcomes. ⑤ Feedback Mechanisms: Feedback loops exist within the economy where macro-policy impacts micro-agents, and the micro-agents’ behaviors, in turn, influence policy decisions.
Given the aforementioned key points, Dynamic Stackelberg Mean Field Games (Dynamic SMFGs) are ideal for macroeconomic policy-making due to:
1. Leader-Follower Dynamics (②⑤): The government, acting as the leader, sets policies first. Households, as followers, adjust their behaviors based on these policies, which captures the strategic interactions between policy-setters and economic agents.
2. Mean Field Interaction (①⑤): The collective behavior of numerous micro-agents is efficiently modeled as a "mean field", which simplifies the analysis of large-scale interactions and their overall impact on the macroeconomy.
3. Dynamic Strategy (②③④): The dynamic nature of these games reflects the evolving economic conditions and policy-making, enabling the model to adapt to new information and economic shocks.
3.2 Dynamic Stackelberg Mean Field Games
Following this, we consider a dynamic SMFG over discrete periods $t\in\{0,\ldots,T\}$ for macroeconomic policy-making issues: At each time $t$, the leader first selects an action $a_{t}^{l}\in\mathcal{A}^{l}$ based on the leader’s state $s_{t}^{l}\in\mathcal{S}^{l}$ and policy $\pi^{l}:\mathcal{S}^{l}\rightarrow\mathcal{A}^{l}$. Subsequently, followers simultaneously determine their actions based on the leader’s action $a^{l}_{t}$ and their private states $s^{f}_{t}\in\mathcal{S}^{f}$. The representative follower’s action $a_{t}^{f}\in\mathcal{A}^{f}$ is derived from a shared policy $\pi^{f}:\mathcal{S}^{f}\times\mathcal{A}^{l}\rightarrow\mathcal{A}^{f}$. The sequences $\{\pi_{t}^{l}\}_{t=0}^{T}$ and $\{\pi_{t}^{f}\}_{t=0}^{T}$ are denoted as $\pi^{l}$ and $\pi^{f}$, respectively.
Followers’ Mean Field Game
Given the leader’s policy $\pi^{l}$, the dynamic SMFG is simplified into a mean field game for the followers. At time $t\in\{0,...,T-1\}$, given the leader’s action $a^{l}_{t}$, the representative follower’s state $s^{f}_{t}$ and action $a^{f}_{t}$, and the population state-action distribution of all followers denoted as $L^{f}_{t}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the representative follower receives a reward $r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ and move to the next state $s_{t+1}^{f}\sim\mathbb{P}(\cdot|s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$. Each follower aims to find the optimal policy $\pi^{f}$ that maximizes his cumulative reward over the time horizon:
$J^{f}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{f}\left[\sum_{t=0}^{T}r^{f}(s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})\right]$ (1)
where $\mathbb{E}^{f}$ denotes the expectation with respect to initial state $s_{0}^{f}\in\mu_{0}^{f}$, $(s_{t+1}^{l},s_{t+1}^{f})\sim P(\cdot\mid s_{t}^{f},a_{t}^{f},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a^{l}_{t}\sim\pi^{l}(\cdot\mid s_{t}^{l}),a^{f}_{t}\sim\pi^{f}(\cdot\mid s_{t}^{f},a^{l}_{t})$ for $t\in\{0,...,T\}$.
Definition 3.1.
[Followers’ Best Response for Leader’s Policy] Given any leader’s policy $\pi^{l}\in\Pi^{l}$ and the followers’ state-action distribution $L^{f}\in\mathcal{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$, the followers’ best response policy ${\pi^{f}}^{*}(\pi^{l},L^{f})$ is defined as
${\pi^{f}}^{*}(\pi^{l},L^{f})\in\arg\max_{\pi^{\prime}}J^{f}(\pi^{l},\pi^{\prime},L^{f}).$
Leader’s Game
At time $t\in\{0,...,T-1\}$, given the leader’s state $s^{l}_{t}$, action $a^{l}_{t}$, and the followers’ state-action distribution $L_{t}^{f}$, the leader will receive a reward $r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})$, and move to a new state $s_{t+1}^{l}$ according to the transition probabilities $\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$. The leader aims to find the optimal policy $\pi^{l}$ to maximize the total expected reward:
$J^{l}(\pi^{l},\pi^{f},L^{f})=\mathbb{E}^{l}\left[\sum_{t=1}^{T}r^{l}(s_{t}^{l},a_{t}^{l},L^{f}_{t})\right],$ (2)
where $\mathbb{E}^{l}$ denotes the expectation under this probability distribution for the leader when $s_{0}^{l}\sim\mu_{0}^{l},s_{t+1}^{l}\sim\mathbb{P}(\cdot|s_{t}^{l},a_{t}^{l},L^{f}_{t})$ for $t\in\{0,...,T-1\}$, and $a_{t}^{l}\sim\pi_{t}^{l}(\cdot\mid s_{t}^{l})$ for $t\in\{0,...,T\}$.
Definition 3.2 (Leader’s Optimal Policy in Dynamic Stackelberg Mean Field Games).
Considering the followers’ best response $(\pi^{f},L^{f})$ for the leader’s policy, which satisfies Definition 3.1, the leader’s optimal policy ${\pi^{l}}^{*}$ in dynamic Stackelberg mean field games is defined as follows:
${\pi^{l}}^{*}$ $\in\mathop{\arg\max}_{{\pi^{l}}^{\prime}}J^{l}({\pi^{l}}^{\prime},\pi^{f},L^{f}),$
[…]
where the followers’ state-action distribution $L^{f}$ satisfies the following Mckean-Vlasov equation:
$L^{f}_{t+1}(s_{t+1}^{f},a_{t+1}^{f})=\sum_{s_{t}^{l},a_{t}^{l},s_{t}^{f},a_{t}^{f}}L^{f}_{t}(s_{t}^{f},a_{t}^{f})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})\mu_{t}^{l}(s_{t}^{l})P(s_{t+1}^{f}\mid s_{t}^{f},a_{t}^{f},a_{t}^{l})\pi_{t+1}^{f}(a_{t+1}^{f}\mid s_{t+1}^{f}),$
$\mu_{t+1}^{l}(s_{t+1}^{l})=\sum_{s_{t}^{l},a_{t}^{l}}\mu_{t}^{l}(s_{t}^{l})\pi_{t}^{l}(a_{t}^{l}\mid s_{t}^{l})P(s_{t+1}^{l}\mid s_{t}^{l},a_{t}^{l}).$ (3)
In conclusion, to study the issue of macroeconomic policy-making, we model it as a dynamic Stackelberg mean field game. This modeling approach aims to learn the optimal policy of the leader, namely the social planner, while considering the followers’ best response over discrete timesteps (Definition 3.2).
3.3 Metrics for Dynamic SMFGs
We will utilize the following three metrics to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the followers’ policy from the best response to the leader’s policy; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
Leader’s Payoff
We define the leader’s payoff using the long-term expected rewards of the leader’s policy $\pi^{l}$ over discrete timesteps, as detailed in Equation 2.
Exploitability
Exploitability is a widely used metric for evaluating convergence and measuring the distance between the representative follower’s policy $\pi^{f}$ and the best response ${\pi^{f}}^{*}$ for the leader’s policy $\pi^{l}$. Formally, it quantifies the extent to which a single follower can improve their payoff by deviating from the population’s behavior to adopt a different policy. Given the leader’s policy $\pi^{l}$ and the followers’ state-action distribution $L^{f}$, the exploitability $\mathcal{E}(\pi^{f};\pi^{l})$ of the follower’s policy $\pi^{f}$ is defined by comparing the value functions induced by $\pi^{f}$ and its best response $\pi^{f*}$:
$\mathcal{E}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L^{f})-J^{f}(\pi^{l},{\pi^{f}},L^{f}).$
Social Optimum and Social Welfare
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare [2, 40]. Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions of the $N$ followers in Equation 1 when $N\to\infty$:
$\mathcal{SW}(\pi^{l},\pi^{f})=\sum^{N}J^{f}\left(\pi^{l},\pi^{f},L^{f}\right).$
4 Stackelberg Mean Field Reinforcement Learning
However, solving for the leader’s optimal policy in dynamic SMFGs, as defined in Definition 3.2, presents several challenges, such as vast state and action spaces, strategic dependencies and information asymmetry between the leader and followers, and potentially conflicting objectives including the leader’s payoff, followers’ utilities, and social welfare. These complexities make devising a model-based algorithm with theoretical proof extremely difficult. Previous works have attempted simplifications using the myopic followers setting [63] or linear-quadratic MFGs [6], but this approach fails to align with macroeconomic policy-making issues. Consequently, we have designed the Stackelberg Mean Field Reinforcement Learning (SMFRL) algorithm to address the above limitations. This is a model-free approach that operates independently of environmental transition information.
Figure 2: The architecture of SMFRL algorithm.
Considering the dynamic SMFGs with a leader and $N$ homogeneous follower agents, at any time $t\in\{0,\ldots,T-1\}$, the joint action of the leader and follower agents $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\},a^{fi}_{t}=\pi^{f}(s^{fi}_{t})$ impacts the environment and results in a reward $\mathbf{r}_{t}=\{r^{l}_{t},r^{f1}_{t},...,r^{fN}_{t}\}$ and the next state $\mathbf{s}_{t+1}=\{s^{l}_{t+1},s^{f1}_{t+1},...,s^{fN}_{t+1}\}$. The experience replay buffer $\mathcal{D}$ contains the tuples $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$, recording the experiences of all agents. Specifically, we define the leader’s policy as $\pi_{\theta^{l}}$ parameterized by $\theta^{l}$ (abbreviated as $\pi^{l}$), and a shared policy for the followers denoted as $\pi_{\theta^{f}}$ with parameters $\theta^{f}$ (abbreviated as $\pi^{f}$). The SMFRL algorithm incorporates several settings (shown in Figure 2) to address the aforementioned challenges:
Mean Field Approximation
To mitigate the problem of non-stationarity, we consider a central critic Q function $Q(\mathbf{s}_{t},\mathbf{a}_{t}):\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}\times\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}\to\mathbb{R}$ for each agent. However, due to the high dimensionality of the joint state and action, the standard Q network is infeasible to learn. Thus, the SMFRL algorithm employs a mean field approximation, characterizing the followers’ state-action pairs $(\mathbf{s}^{f}_{t},\mathbf{a}^{f}_{t})$ by using the followers’ population distribution $L^{f}_{t}$, which can be presented through various metrics such as mean, variance, probability density functions, skewness, and kurtosis. The theoretical proof is detailed in Appendix F. This approach yields $Q^{l}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$ and $Q^{f}(s^{f}_{t},a^{l}_{t},a^{f}_{t},L^{f}_{t})$ for evaluating the leader and follower agents respectively. The mean field approximation effectively reduces the dimensionality of the inputs to the Q function, thereby facilitating the learning of the critic Q network and ensuring the Q function can evaluate leader and follower policies based on the overall follower distribution.
Leader’s Update for Optimal Policy
In SMFRL, the leader’s policy $\pi_{\theta^{l}}$, is optimized using a deterministic policy gradient approach [54]. The policy gradient for the leader can be estimated as:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right],$
where ${Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$, the leader’s action-value function, is approximated by a neural network parameterized by $\phi^{l}$. This network is periodically updated to minimize the loss:
$\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$
where the target value $y^{l}_{t}$ is given by:
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})},$
and $\phi^{l}_{-}$ represents the parameters of the target Q network, $L^{f}_{t+1}$ is determined based on Mckean-Vlasov equation (3), and $\gamma$ is the discount factor. Differentiating the loss function $\mathcal{L}(\phi^{l})$ yields the gradient utilized for training:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right].$
Followers’ Update for Best Response
To enable follower agents to effectively learn the best response policy $\pi^{f*}$ to the leader’s policy $\pi^{l}$, we employ multiple updates for followers’ shared policy and Q networks in the SMFRL algorithm. Each update cycle refines the policy network $\pi_{\theta^{f}}$ for the followers using the policy gradient method:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f}_{t})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]},$
where the action-value function ${Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})$, parameterized by $\phi^{f}$, is periodically updated by minimizing the mean squared error (MSE) loss:
$\mathcal{L}(\phi^{f})=$ $\mathbb{E}_{{s}_{t}^{f},{a}_{t}^{f},a^{l}_{t},L^{f}_{t},s_{t+1}^{f}\sim\mathcal{D}}\left[\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)^{2}\right],$ (4)
$y^{f}_{t}=$ $r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}),$
with the next leader’s action $a^{l}_{t+1}=\mu_{\phi^{l}_{-}}(s^{l}_{t+1})$, and the next population distribution $L^{f}_{t+1}=L(\mathbf{s}_{t+1}^{f},\mathbf{a}_{t+1}^{f})$. The joint actions of followers $\mathbf{a}_{t+1}^{f}=\{a^{f1}_{t+1},\ldots,a^{fN}_{t+1}\}$ are determined with each action $a^{fi}_{t+1}=\mu(s^{fi}_{t+1},a^{l}_{t+1})$. The gradient of the loss function $\mathcal{L}(\phi^{f})$ is derived as:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
The pseudocode for the SMFRL algorithm (1) is presented in Appendix A.
Table 1: Evaluation of SMFRL and five baselines on 3 metrics for dynamic Stackelberg mean field games with different followers’ numbers.
Metrics Followers’ Optimal SMFRL MF-MARL MADDPG IDDPG DDPG U.S. Policy
Number Value +Real +Real
Leader’s Payoff N=100 \ 3387.36 -76.82 -282.62 -701.61 2840.07 1718.08
N=1000 \ 3411.16 -668.03 121.18 -767.53 2874.19 -796.4
Exploitability N=100 0. 0.000107 0.002190 0.007908 0.001333 0.009885 0.032490
N=1000 0. 0.000158 0.003481 0.035613 0.000565 0.003002 0.010522
Social Welfare N=100 100 96.87 82.02 55.74 83.09 94.5 79.23
N=1000 1000 968.94 834.93 645.38 874.53 943.82 334.79
5 Experiment
In this section, we conduct experiments within the TaxAI environment to answer the following three questions: (1) How effective is our SMFRL algorithm in solving dynamic SMFGs? (2) For macroeconomic policy-making issues, how does our method compare to real-world strategies, economic models, and existing AI-based methods? (3) Given the likely real-world scenario where some households do not adopt the SMFG household policy learned by our method, does our approach remain effective?
TaxAI Environment
TaxAI [41] is one of the rare reinforcement learning environments designed to study optimal macroeconomic policies. It simulates dynamic interactions between large-scale households and the government, making it particularly well-suited for experimenting with SMFGs. TaxAI models the optimal tax policy problems where the government optimizes tax policies to promote economic growth (i.e. GDP) and social equity (i.e. Gini). In response, households optimize their labor supply and savings strategies to maximize their long-term utility.
5.1 Solving Dynamic SMFGs
Metrics and Baselines
We compare several classic multi-agent reinforcement learning (MARL) algorithms based on three metrics: leader’s payoff, exploitability, and social welfare, as defined in Section 3.3. Optimal Value: The ideal value of exploitability is 00, indicating that the follower’s policy achieves the best response to the leader’s policy. The upper bound of Social Welfare is $N$, given that household utility in TaxAI cannot exceed 1. For Leader’s Payoff, no maximum has been established. MF-MARL [61]: Utilizes the DDPG [36] algorithm for the leader and a mean-field MARL approach for the followers. MADDPG [38] and IDDPG [17]: Both feature shared actor and critic networks among followers. DDPG+Real: DDPG for the leader and behavior cloning [57] from real data for followers. U.S. policy+Real: U.S. federal tax policy for the leader with behavior cloning for the followers.
Results Analysis
Table 1 presents the performance metrics for SMFRL and five baselines, alongside their optimal values, averaged over 10 evaluations of models trained with at least five seeds. SMFRL notably outperforms the baselines in achieving higher leader’s payoff, demonstrating its effectiveness in optimizing the leader’s objective. The DDPG+Real method, using a real-data trained policy, attains a substantial payoff but exhibits a significant gap from the households’ optimal response, as indicated by its higher exploitability. In contrast, SMFRL significantly reduces exploitability for both $N=100$ and $N=1000$, suggesting that its policy for households closely approximates the best response to the leader’s strategy, which is crucial for solving Stackelberg mean field games. Moreover, SMFRL closely approaches the optimal social welfare value, indicating that its solutions are the most socially optimal among the considered baselines. In summary, the SMFRL algorithm effectively learns policies that maximize the leader’s objective while ensuring that followers attain the best response to the leader’s policy, closely approaching the social optimum. The training curves are presented in Figure 4, and the test results for macroeconomic indicators are detailed in Table 1 in Appendix D.
5.2 Macroeconomic Policy-Making
Table 2: Test results 6 tax policies across five macroeconomic indicators at two numbers of followers, $N=100$ and $N=1000$.
Households’ Policies Macroeconomic Indicators
Number Per Capita GDP Social Welfare Income Gini Wealth Gini Years
$\mathbf{N=100}$ Free Market 1.37e+05 32.97 0.89 0.92 1.10
Saez Tax 2.34e+12 73.82 0.21 0.38 300.00
US Federal Tax 4.88e+11 94.19 0.40 0.40 289.55
AI Economist-BC 4.24e+12 97.24 0.54 0.52 299.55
AI Economist 1.26e+05 72.81 0.88 0.91 1.00
SMFG 9.59e+12 96.87 0.52 0.51 300.0
$\mathbf{N=1000}$ Free Market 1.41e+05 334.79 0.90 0.93 1.00
Saez Tax 6.35e+11 498.88 0.68 0.73 100.578
US Federal Tax 1.41e+05 351.17 0.89 0.93 1.00
AI Economist-BC N/A N/A N/A N/A N/A
AI Economist N/A N/A N/A N/A N/A
SMFG 1.10e+13 968.94 0.54 0.53 300.00
Metrics and Baselines
In addressing macroeconomic policymaking, we utilize classical macroeconomic indicators from the real world to evaluate the performance of different policies. Per Capita GDP indicates the level of economic development, while Income Gini and Wealth Gini measure the inequality in households’ income and wealth, respectively; a lower Gini index suggests greater social equality. The Years metric represents the sustainable duration of an economy, capped at 300 years; economies may terminate earlier due to excessive inequality or inadequate productivity to meet the populace’s needs. We compare several methods for solving macroeconomic policies: the Free Market [5] policy entails no government intervention; the Saez Tax [50] represents a well-known economic approach, often advocated for specific tax reforms in practice. The U.S. Federal Tax is the real policy from the OECD data  [^2] https://data.oecd.org/. In these approaches, households employ strategies based on behavior cloning from real data. Additionally, the AI Economist [62] represents an AI-based policy using a two-level Proximal Policy Optimization (PPO) framework. The AI Economist-BC method leverages behavior cloning policy on real-world data for pretraining.
Results Analysis
Table 2 presents the average values of macroeconomic indicators across multiple test rounds for different algorithms under varying numbers of households. The results reveal that the Saez Tax method achieves the lowest income Gini and wealth Gini when the number of households is small ($N=100$). However, our SMFG method yields higher per capita GDP and social welfare. At $N=1000$, the advantages of the SMFG algorithm become even more pronounced compared to other methods. It secures the longest duration (Years), the highest GDP, and the lowest income Gini and wealth Gini. The social welfare derived from the SMFG approach closely approaches the optimal value of 1000. This demonstrates that our method can learn the macroeconomic policy that simultaneously optimizes economic growth, social equity, and social welfare, surpassing existing economic approaches like the Saez Tax and AI-based policies (i.e. AI Economist). We provide the corresponding training curves in Figure 5 in Appendix D.
5.3 Effectiveness Analysis in Diverse Households
Figure 3: The left subfigure illustrates a scenario with heterogeneous households adopting various policies. The right subfigure presents test results for households’ decisions (consumption, labor), microeconomic indicators (utility, wealth, income), and macroeconomic indicators (GDP) when there are different proportions of SMFG households ($N=100$).
Problem Description
The above experiments have already demonstrated the effectiveness of our method in addressing macroeconomic policy-making problems. However, when extending SMFG approach to real-world scenarios, where governments and households implement leader and follower policies trained via the SMFG method (referred to as SMFG government and SMFG households), there remain partial households that might not adopt the SMFG follower policy. In economic terms, this scenario can be interpreted as the government introducing new policies and providing recommendations to households, some of whom may choose not to follow these recommendations. When faced with heterogeneous households that adopt varying strategies, the question arises: is our SMFG method still effective? In experiments, to model households that do not adopt the SMFG policy, referred to as Real-data Households, we utilize behavior cloning policy trained by real data. Related experimental details are shown in Appendix D.4. We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators. In Figure 3, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis). GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
Result Analysis
From Figure 3, we identify several key findings: (1) SMFG Households outperform Real-data Households across microeconomic indicators such as wealth, income, and utility, which should encourage more households to adopt the SMFG policy; (2) The presence of heterogeneous households does not prevent SMFG Households from achieving near-optimal utility, capped at 1 in the TaxAI environment; (3) The lower labor contribution from Real-data Households leads to a decrease in Per Capita GDP as their proportion increases. Thus, the more households adopt the SMFG follower policy, the more effective the SMFG method becomes.
6 Conclusion
This paper innovatively models macroeconomic policymaking as a dynamic Stackelberg mean field game, with the government acting as the leader and large-scale households as followers. We introduce a data-driven Stackelberg mean field reinforcement learning algorithm that effectively solves SMFGs. When applied to macroeconomic policy formulation, our method significantly outperforms existing economic approaches and AI-based policies, especially in scenarios involving large numbers of households. Furthermore, we demonstrate the robustness of our approach, showing that it retains effectiveness even when some households do not adopt the SMFG policy. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
Appendix A SMFRL algorithm pseudocode
Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
  Initialize ${Q}_{\phi^{l}},{Q}_{\phi^{l}_{-}},{Q}_{\phi^{f}},{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
  for epoch = 1 to M do
     Receive initial state $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
     for  $t=1$ to max-epoch-length do
        For the leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t. the current policy and exploration.
        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next state $\mathbf{s}_{t+1}$.
        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
        for j =1 to update-cycles do
           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
           For each follower agent, perform inner updates:
           for k = 1 to inner-update-cycles do
              Set
$y^{f}_{t}=r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
              Update follower’s critic by minimizing the loss:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
              Update follower’s actor by policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
           end for
           After follower updates, update the leader’s networks:
           Set
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
           Update leader’s critic net by minimizing the loss:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
           Update leader’s actor by sampled policy gradient:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
        end for
        Update target network parameters:
$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
     end for
  end for
Appendix B Assumptions and Limitations
Assumptions
This paper models the problem of macroeconomic policy-making as a Dynamic Stackelberg Mean Field Game, based on the following assumptions: (1) Homogeneous followers: We assume that a large-scale group of households is homogeneous. They can use different characteristics as observations to influence decisions, but there are commonalities in human behavioral strategies. (2) Rational Expectations: We assume that both macro and micro agents engage in rational decision-making, adjusting their future expectations based on observed information. However, in reality, the level of rationality varies among different households. Most households exhibit bounded rationality, and their expectations and preferences differ accordingly. (3) Experimental environment: We validate our approach through experiments in the TaxAI environment, based on the assumption that results within this environment can provide insights applicable to real-world scenarios. Addressing and potentially relaxing these assumptions will be a primary focus of our future research.
Limitations
The limitations of our SMFG method will be thoroughly investigated in future work: (1) We plan to consider dynamic games involving multiple leaders and large-scale followers to explore policy coordination across various macroeconomic sectors. (2) We will continue to develop theoretical proofs for the equilibrium solutions in Stackelberg mean field games. Currently, our approach is empirically demonstrated by showing that followers converge toward their best responses and that the leader achieves higher performance compared to other baselines. (3) We intend to examine dynamic games between a leader and a large, heterogeneous group of followers, including scenarios where followers dynamically alter their strategies, to determine the optimal leader policy. Addressing these limitations will provide further insights applicable to real-world scenarios.
Appendix C Impact Statements
This paper presents work whose goal is to advance the fields of AI for social science and AI for economics. Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix D Additional Results
Compute Resources
All experiments are run on 2 workstations: A 64-bit server with dual AMD EPYC 7742 64-Core Processors @2.25 GHz, 256 cores, 512 threads, 503GB RAM, and 2 NVIDIA A100-PCIE-40GB GPU. A 64-bit workstation with Intel Core i9-10920X CPU @ 3.50GHz, 24 cores, 48 threads, 125 GB RAM, and 2 NVIDIA RTX2080 Ti GPUs. The following Table 3 shows the approximate training times for several algorithms.
Table 3: the approximate training times for several algorithms for solving dynamic SMFGs.
Algorithm N=100 (hours) N=1000 (hours)
SMFRL 4 14
MADDPG 3 16
MF-MARL 4 9
IDDPG 2 6
DDPG+Real 1 5
Rule-based+Real 0.25 2
D.1 Training Curves for Dynamic SMFGs
Due to the challenging nature of the task, the baselines we compared did not perform well in solving dynamic SMFGs. Therefore, we utilized behavior cloning as a pre-training strategy for the follower agents. Therefore, we compared SMFRL with 5 baselines in 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 4 illustrates the training curves of 4 key macroeconomic indicators under these three settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
Figure 4 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, SMFRL, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The SMFRL algorithm substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, SMFRL achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, SMFRL consistently obtains the most optimal solutions across all indicators.
Table 4: Performance metrics of six algorithms across different settings on key economic indicators.
Settings Algorithm GDP per Capita Social Welfare Income Gini Wealth Gini Years
N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
MADDPG 4.93e+04 55.74 0.91 0.92 1.00
IDDPG 1.21e+05 83.09 0.88 0.92 1.00
MF-MARL 8.66e+07 82.02 0.82 0.83 75.75
SMFRL 9.59e+12 96.87 0.52 0.51 300.00
N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
MADDPG 3.37e+09 645.38 0.83 0.85 63.58
IDDPG 1.33e+05 874.53 0.89 0.92 1.00
MF-MARL 1.31e+05 834.93 0.88 0.92 1.50
SMFRL 1.10e+13 968.94 0.54 0.53 300.00
N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
MADDPG 6.38e+12 93.89 0.57 0.58 268.53
IDDPG 7.41e+12 98.16 0.53 0.55 300.00
MF-MARL 5.44e+12 98.21 0.50 0.52 300.00
SMFRL 1.01e+13 96.90 0.51 0.53 299.89
N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
MADDPG 6.82e+12 954.88 0.56 0.62 278.50
IDDPG 2.79e+12 512.19 0.77 0.81 100.68
MF-MARL 1.13e+05 440.00 0.90 0.93 1.00
SMFRL 9.68e+12 975.15 0.52 0.51 300.00
Figure 4: The training curves for 6 algorithms on 4 macroeconomic indicators, comparing settings without behavior cloning as pre-train (N=100 & N=1000) and with behavior cloning (N=100-BC & N=1000-BC).
D.2 Training Curves for Various Tax Policies
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 5 displays the training curves and Table 2 shows the test results in Appendix. Both Figure 5 and Table 2 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
Figure 5: The training curves for 6 tax policies on 5 macroeconomic indicators (N=100 & N=1000 with BC)
D.3 Efficiency-Equity Tradeoff of Policies
Figure 6: Comparative performance of various policies under multi-objective assessment (Efficiency-Equity).
In economics, the Efficiency-Equity Tradeoff is a highly debated issue. We find that our SMFG method is optimal in balancing efficiency-equity, except in cases of extreme concern for social fairness. In our study, we depict the economic efficiency (Per capita GDP) on the Y-axis and equity (wealth Gini) on the X-axis of Figure 6(a) for various policies. Different policies are represented by circles of different colors, with their sizes proportional to social welfare. Different circles of the same color correspond to different seeds. Figure 6 (a) shows that the wealth Gini indices for SMFG and AI Economist-BC are similar, but SMFG has a higher GDP, suggesting its superiority over AI Economist-BC. SMFG significantly outperforms the free market policy and AI Economist due to its higher GDPs and lower wealth Ginis. However, comparing SMFG with the Saez tax and the U.S. Federal tax policy in terms of both economic efficiency (GDP) and social equity (Gini) is challenging. Therefore, we introduce Figure 6 (b) to demonstrate the performance of different policies under various weights in a multi-objective assessment.
In Figure 6 (b), the Y-axis shows the weighted values of the multi-objective function $Y=log(\text{per capita GDP})+\alpha(\text{wealth Gini})$, and the X-axis represents the weight of the wealth Gini index. For each weight $\alpha$, we compute the multi-objective weighted values for those policies, represented as circles of different colors. Due to the logarithmic treatment of GDP in (b), when $\alpha=10$, the overall objective focuses solely on social fairness; when $\alpha=0$, the overall objective is concerned only with efficiency. Our findings in (b) reveal that only when $\alpha\geq 8$, which indicates a substantial emphasis on social equity, does the Saez tax outperform SMFG. However, SMFG consistently proves to be the most effective policy under a wide range of preference settings.
D.4 Behavior Cloning Experiments
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.3 to compare with SMFG Households. We use the statistical data from the 2022 Survey of Consumer Finances (SCF)  [^3] https://www.federalreserve.gov/econres/scfindex.htm. Details about the behavior cloning experiments are provided below. This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right). The loss convergence curve of behavior cloning is shown in Figure 7. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and MADDPG can converge to below 0.1.
Figure 7: The behavior cloning loss for 4 networks in two loss types.
Appendix E Saez tax
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows [50]. The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
$\tau(z)=\frac{1-G(z)}{1-G(z)+\alpha(z)e(z)}$
To further elaborate, the marginal average income at a given income level $z$, normalized by the fraction of incomes above $z$, is denoted as $\alpha(z)$.
$\alpha(z)=\frac{zf(z)}{1-F(z)}$
The reverse cumulative Pareto weight over incomes above $z$ is represented by $G(z)$.
$G(z)=\frac{1}{1-F(z)}\int_{z^{\prime}=z}^{\infty}p\left(z^{\prime}\right)g\left(z^{\prime}\right)\mathrm{d}z^{\prime}$
From the above calculation formula, we can calculate $G(z)$ and $\alpha(z)$ by income distribution. We obtain the data of income and marginal tax rate through the interaction between the agent and environment and store them in the buffer. It is worth noting that the amount of buffer is fixed.
To simplify the environment, we discretize the continuous income distribution, by dividing income into several brackets and calculating a marginal tax rate $\tau(z)$ for each income range. Within each tax bracket, we determine the tax rate for that bracket by averaging the income ranges in that bracket. In other words, income levels falling within the income range are calculated as the average of that range. In particular, when calculating the top bracket rate, it is not convenient to calculate the average because its upper limit is infinite. So here $G(z)$ represents the total social welfare weight of incomes in the top bracket, when calculating $\alpha(z)$, we take the average income of the top income bracket as the average of the interval.
Elasticity $e(z)$ shows the sensitivity of the agent’s income z to changes in tax rates. Estimating elasticity is very difficult in the process of calculating tax rates, here we estimate the elasticity $e(z)$ using a regression method through income and marginal tax rates under varying fixed flat-tax systems, which produces an estimate equal to approximately 1.
$e(z)=\frac{1-\tau(z)}{z}\frac{dz}{d(1-\tau(z))}$
$\log(Z)=\hat{e}\cdot\log(1-\tau)+\log\left(\widehat{Z}^{0}\right)$
where $Z=\sum_{i}z_{i}$ when tax rates is $\tau$.
Appendix F Theoretical Proof
Proposition F.1.
The action value functions $Q_{i}$ and ${Q}_{t}$ are equivalent, i.e. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})={Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
Proof.
We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$. For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$. For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$). Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form. We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions. The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$. The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})={Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$. ∎
Appendix G Economic Model Details
Economic activities among households aggregate into labor markets, capital markets, goods markets, etc. In the labor market, households are the providers of labor, with the aggregate supply $S(W_{t})=\sum_{i}^{N}e_{t}^{i}h_{t}^{i}$, and firms are the demanders of labor, with the aggregate demand $D(W_{t})=\mathcal{L}_{t}$. When supply equals demand in the labor market, there exists an equilibrium price $W_{t}^{*}$ that satisfies:
$S(W^{*}_{t})=D(W^{*}_{t}),\mathcal{L}_{t}=\sum^{N}_{i}e_{t}^{i}h_{t}^{i}.$
In the capital market, financial intermediaries play a crucial role, lending the total deposits of households $A_{t+1}=\sum_{i}^{N}a_{t+1}$ to firms as production capital $K_{t+1}$, and purchasing government bonds $B_{t+1}$ at the interest rate $r_{t}$. The capital market clears when supply equals demand:
$K_{t+1}+B_{t+1}-A_{t+1}=\left(r_{t}+1\right)(K_{t}+B_{t}-A_{t})$
In the goods market, firms produce and supply goods, while all households, the government, and physical capital investments $X_{t}$ demand them. The goods market clears when:
$Y_{t}=C_{t}+G_{t}+X_{t}$
where $C_{t}=\sum_{i}^{N}c^{i}_{t}$ represents the total consumption of consumers, and $G_{t}$ is government spending. The supply, demand, and price represent the states of the market.